Subjective wellbeing was measured among a sample of Statistics students with M=150 and S squared=25. Determine the z-scores for the students who obtained the following scores on the subjective wellbeing measure.

a) 110

b) 135



1
Expert's answer
2021-05-31T16:13:02-0400

"\mu = 150 \\

S^2=25 \\

S=5 \\

Z=\frac{X- \mu}{S}"

a)

"Z=\frac{110-150}{5}=-8"

b)

"Z = \frac{135-150}{5}=-3"